Explain the difference between the required rate of return and the expected rate of return. When comparing the required rate of return to the expected rate of return, when would we purchase the stock (i.e., when the required rate is <,>, or = to the expected rate)?
Answer #1

Expected Return:

              Expected return is the rate of return which firm expects from investment. It represents return from the stock. For example if price of a stock is 100 today and same will fetch 112 next year then expected return is said to be 12 %. It is calculated as follows : (112-100)-1 = 1.12-1 = 0.12 = 12%

Required Rate of Return:

                    Required Rate of Return is the minimum rate of return the firm has to earn. It depends on the risk of the bond i.e. certainty of expected return future benefits from the bond. If risk is higher the required return will be high and vice versa.

Comparing required rate of return and the expected rate of return:

                                                                Expected return is what the stock earns and required rate of return is what the investor expects. The investor takes investment decision after comparing required return and expected return. If expected return (return the investment earns) is greater than his required return, then he will invest in that investment.

Condition

Decision

Expected Return > Required Return

Invest

Expected Return < Required Return

Not to Invest

Expected Return = Required Return

Invest